In the preparation of an intravenous infusion, a vial containing 115 mg of the drug is diluted to 5 mL with sodium chloride for injection. Then, the contents of the vial are added to 110 mL of an infusion solution. Calculate the drug strength of the final infusion in (a) mg/mL, (b) percent strength, and (c) ratio strength.
Answer #1

(a) Drug dosage is 115 mg

Diluted with 5 ml sodium chloride + 110 ml infusion solution = 115 ml

115 mg/ 115 ml = 1 mg/ 1 ml

(b) Percent strength weight in volume w/v

This is measured by calculating drug (grams) in 100 ml of solution

115 mg = 0.115 gm (above given solution)

Percent strength = weight of solvent \div volume of solution \times 100

                           = 0.115 \div 115 \times 100

                           = 0.1 % w/v

(c) Ratio strength is c1 \div v1 = c2 \div v2

               0.1 \div 115 = 1 \div x

             x = 115 \times 1 \div 0.1

                = 1150

The ratio strength is 1 part in 1150 that is 1:1150